# TabArena: A Living Benchmark for Machine Learning on Tabular Data

### v2

@@ Abstract
With the growing popularity of deep learning and foundation models for tabular data, the need for standardized and reliable benchmarks is higher than ever. However, current benchmarks are static. Their design is not updated even if flaws are discovered, model versions are updated, or new models are released. To address this, we introduce TabArena, the first continuously maintained living tabular benchmarking system. To launch TabArena, we manually curate a representative collection of datasets and well-implemented models, conduct a large-scale benchmarking study to initialize a public leaderboard, and assemble a team of experienced maintainers. Our results highlight the influence of validation method and ensembling of hyperparameter configurations to benchmark models at their full potential. While gradient-boosted trees are still strong contenders on practical tabular datasets, we observe that deep learning methods have caught up under larger time budgets with ensembling. At the same time, foundation models excel on smaller datasets. Finally, we show that ensembles across models advance the state-of-the-art in tabular machine learning and investigate the contributions of individual models. We launch TabArena with a public leaderboard, reproducible code, and maintenance protocols to create a living benchmark available at https://tabarena.ai.
Figure 1: TabArena-v0.1 Leaderboard. We evaluate models under default parameters, tuning, and weighted ensembling [1] of hyperparameters. Since TabICL and TabPFNv2 are not applicable to all datasets, we evaluate them on subsets of the benchmark in Figure 4 .
@@ 1 Introduction
[…]
@@
TabArena-v0.1 Focus Statement. TabArena focuses on evaluating predictive machine learning models for tabular data. Our long-term vision is to make TabArena representative for all use cases of tabular data. For TabArena-v0.1, we initialize the benchmark focusing on the most common type of tabular machine learning problem: Tabular classification and regression for independent and identically distributed (IID) data, spanning the small to medium data regime. We explicitly leave for future work use cases such as non-IID data (e.g., temporal dependencies, subject groups, or distribution shifts); few-shot predictions, or very small data (e.g., less than 500 training samples); large data (e.g., more than 250 000250000250\,000250 000 training samples); or other machine learning tasks such as clustering, anomaly detection, subgroup discovery, or survival analysis.
@@ 2 TabArena
TabArena is a living benchmark because of our protocols, which govern the curation of: (2.1) models and hyperparameter optimization; (2.2) datasets; (2.3) and evaluation design. Through continuous application and refinement of these protocols, we ensure TabArena remains current and maintained.
@@ 2.1 Models and Hyperparameter Optimization Protocol
TabArena models are powered by three components: (1) implementation in a well-tested modeling framework used in real-world applications; (2) curated hyperparameter optimization protocols; (3) improved validation and ensembling strategies, including ensembling over instances of a single model class. Table 1 provides an overview of the models benchmarked in TabArena-v0.1.
@@
Linear / Logistic Regression Linear TabRepo [Uncaptioned image]
K-Nearest Neighbors KNN TabRepo [Uncaptioned image]
Implementation Framework.  For implementing models, we rely on functionalities from AutoGluon [19], an established machine learning framework used in practical applications. Each model is implemented within the standardized AbstractModel framework, which aligns with the scikit-learn [24] API, and includes: (1) model-agnostic preprocessing, (2) support for (inner) cross-validation with ensembling, (3) hyperparameter optimization, (4) evaluation metrics, (5) fold-wise training parallelization, (6) (customizable) model-specific preprocessing pipeline, (7) (customizable) early stopping and validation logic, and (8) unit tests. As a result, any model implemented in TabArena can be readily deployed for real-world use cases or within predictive machine learning systems. Moreover, the pipeline logic encompassing models within TabArena is implemented in a tested framework regularly used in real-world applications. Section C.1 summarizes further implementation details, and Section D.2 includes a detailed protocol for contributing models.
Cross-validation and Ensembles. As can be seen in various Kaggle competitions and academic studies, cf. [25, 8, 26, 20], for most datasets, peak performance requires ensembling strategies. Therefore, we default to using $8$-fold cross-validation ensembles [27] within TabArena. For all foundation models, we refit on training and validation data instead of using cross-validation ensembles, following recommendations from the authors of TabPFN and TabICL. In addition, we evaluate each tunable model using post-hoc ensembling [1] of different hyperparameter configurations.
Hyperparameter Optimization.  For each model, we curate a strong hyperparameter search space; for full details, see Appendix C.2. Where possible, we started with the search spaces from the original paper and finalized them in dialogue with the models’ authors. Otherwise, we curated search spaces from prior work. We evaluate $1$ default and a fixed set of $200$ randomly-sampled hyperparameter configurations for all models, except for TabICL and TabDPT. TabICL and TabDPT do not specify hyperparameter optimization (HPO) per the original paper and implementation; thus, we restrict ourselves to evaluating only their in-context learning performance. Each hyperparameter configuration is validated using $8$-fold (inner) cross-validation. We use the score of this (inner) cross-validation to select the best hyperparameter configuration.
For practical reasons, we restrict the time to evaluate one configuration on one train split of a dataset to $1$ hour. We analyze the impact of this time limit in Section A.2.
@@ 2.2 Datasets Protocol
Table 2: Comparison of Tabular Benchmarks. We systematically compare prior tabular benchmark studies across six characteristics. Inner and outer splits: the number of splits used for inner or outer validation: $1$ for holdout validation; - if the benchmark does not specify; and any other number specifies the total number of splits from (repeated) cross-validation. If a set is given, the benchmark uses different splits for different datasets. Ensembling: Whether the benchmark studies ensembling of configurations for individual models (), uses any other ensembling ((✓)), or uses no ensembling at all (). Manual Curation: Whether the benchmark filters datasets based on criteria beyond simple automation, () or not (). Datasets remaining: the number of datasets remaining after filtering by our criteria. Results available: Whether the benchmark shares no re-usable results (), only metric results ((✓)), or metric results and predictions (). HPO Limit: How hyperparameter optimization was limited in the number of configurations and/or hours.
@@
Many existing benchmarks were curated using semi-automated procedures to collect datasets according to simple characteristics, often to obtain as many datasets as possible. In contrast, we reject the notion of automatically collecting datasets without any sanity check. Instead, we focus on carefully, manually curating a representative collection of datasets. Although some previous benchmarks manually curated data, most of the included datasets do not meet the criteria of IID tabular datasets, as seen in Table 2. Our work indicates a turning point with the most extensive and conscientious manual curation effort for machine learning on IID tabular data so far.
We define criteria for data selection according to our focus statement and filter 1053 datasets used in 13 prior benchmarks accordingly. Figure 2 describes our selection process. Notably, only the deduplication step and size filters can be automated. Faithfully applying the other criteria requires manual human effort per-dataset, demonstrating the downsides of automated data curation procedures.
@@
Dataset Selection Criteria.  We comprehensively describe our selection criteria in Section B.1. In short, we selected datasets that fulfilled the following requirements: (1) The dataset and its predictive machine learning task are unique within our benchmark; (2) The dataset is IID, that is, a random split is appropriate for the underlying original task; (3) The dataset is not from a non-tabular modality, such as images, where it is unclear whether tabular machine learning is a reasonable alternative to domain-specific methods; (4) The dataset stems from a real random distribution, and is not generated, e.g., from a deterministic function; (5) The dataset was published explicitly for a predictive modeling task in a real-world application; (6) The dataset is small-to-medium-sized, i.e., it has at least $500$ and at most 250 000250000250\,000250 000 train samples; (7) We can use a version of the dataset without pre-applied problematic preprocessing, such as irreversible data leaks; (8) The dataset was originally published with a license allowing for scientific usage; (9) The dataset and its structured metadata can be automatically downloaded via a public API, or we are allowed to upload the dataset to a public API; (10) The dataset and its predictive task do not raise (obvious) ethical concerns.
@@
Figure 3 summarizes the characteristics of our datasets; we share per-dataset details in Section B.2. Throughout our curation process, we noticed, among others, two trends. First, across benchmarks, the versions of datasets were often inconsistent. Benchmarks used the same name and source for a dataset, but with different preprocessing, features, or even targets. Thus, dataset-specific performance comparisons across benchmarks were often invalid. Second, the number of datasets that are truly suitable for benchmarking IID tabular data approaches is surprisingly small after carefully inspecting the tasks represented by datasets. We share more noteworthy observations in Section B.3.
Call for Data Contributions.TabArena is a living benchmark, and the datasets we curated are a significant part of this living, continuously maintained system. As mentioned above, we invite the community to scrutinize our curation efforts. Moreover, to keep TabArena up-to-date, we implore the broader scientific community to share their tabular data publicly. We are also actively looking for contributions of data outside of TabArena-v0.1’s focus, i.e., data that is non-IID, tiny, or large, as well as tasks where extracing features from other modalities is still considered a reasonable approach. Likewise, we invite the community to curate additional existing datasets. We provide a detailed protocol to contribute a new dataset to TabArena in Section D.3.
@@ 2.3 Evaluation Design Protocol
The TabArena leaderboard aims to assess the state-of-the-art for predictive machine learning. With this in mind, we design the evaluation to produce a reliable and representative leaderboard. To guard against randomness in the data and method significantly impacting our conclusions, we repeat our experiments per dataset. We employ a dataset-specific repetition strategy using more repeats for smaller datasets and fewer for larger datasets, because the significance of randomness decreases with dataset size. The strategy is as follows: (I) for datasets with less than 2 50025002\,5002 500 samples, we use $10$ times repeated $3$-fold outer cross-validation; (II) for all other datasets, we use $3$ repeats.
We evaluate models using the Elo rating system [38], following Chatbot Arena [39]. Elo is a pairwise comparison-based rating system where each model’s rating predicts its expected win probability against others, with a 400-point Elo gap corresponding to a 10 to 1 (91%) expected win rate. We calibrate 1000 Elo to the performance of our default random forest configuration across all figures, and perform 100 rounds of bootstrapping to obtain 95% confidence intervals, as in [39]. In our main results, Elo scores are computed using ROC AUC for binary classification, log-loss for multiclass classification, and RMSE for regression. The leaderboard offers just one view of the results; we also track additional aspects, such as alternative metrics, training-, and inference-efficiency, to enable diverse analyses. Furthermore, we provide scripts as part of TabArena to generate the leaderboard, evaluation plots, and inspect the results.
TabArena Reference Pipeline. Following the recommendations of [11], we include a reference pipeline in our benchmark. This reference pipeline is applied independently of the tuning protocol and constraints we constructed for models within TabArena. The reference pipeline aims to represent the performance easily achievable by a practitioner on a dataset.
@@
Living Benchmark.TabArena marks the start of an open-source initiative towards a continuously updated, collectively shared assessment of the state-of-the-art in tabular machine learning. Therefore, as one of the most crucial parts of TabArena, we define a protocol for researchers and practitioners to submit models to the live leaderboard. The protocol is detailed in Section D.4
@@ 3 Results
To initialize the leaderboard, we run $16$ curated models (Section 2.1) on $51$ curated datasets (Section 2.2) within TabArena’s evaluation design (Section 2.3). Likewise, we evaluate AutoGluon, the reference pipeline, on all datasets. We run TabM, ModernNCA, and the foundation models on GPU and all other models, as well as AutoGluon, on CPU. We give all runs $32$ GB of RAM and $100$ GB of disk space. We perform CPU runs in the cloud via Amazon Web Services on an M6i.2xlarge EC2 instance (eight cores Intel Xeon CPU). We perform GPU runs on an NVIDIA L40S with 48 GB VRAM and eight cores of an AMD EPYC 9334 CPU. Based on our logs, the compute time for all our experiments without parallelization, including overhead such as scheduling, is ${\sim}15$ wall-clock years.
@@
Figure 1 shows our main results, the leaderboard of TabArena-v0.1, which includes the performance of methods with default parameters, hyperparameter tuning, and weighted post-hoc ensembling [1] of different hyperparameter configurations. Furthermore, we present one leaderboard encompassing all models in Section A.1, where we impute missing datasets, representing the live leaderboard available on tabarena.ai. To provide a wide range of perspectives, we also share additional versions of the leaderboard in Section A.1, such as using alternative metrics or task-wise dataset subsets. Finally, we analyze the statistical significance of our comparison in Section A.5.
@@
Post-Hoc Ensembled Deep Learning Models Dominate the Leaderboard. In line with previous work [33], CatBoost is ranked first in the conventional tuning regime (Figure 1). However, the TabArena-v0.1 leaderboard reveals that after post-hoc ensembling, neural networks are the strongest single models on average in TabArena-v0.1. Our results show that the peak performance of models is misrepresented unless post-hoc ensembling is used. This is evident in the observation that the top three models in our leaderboard (TabM, LightGBM, RealMLP; see Figure 1) would all be worse than the actual fourth-best model (CatBoost) without post-hoc ensembling. While practitioners must trade off the increased predictive performance with the increased inference cost (see Figure 5) [40], peak performance requires post-hoc ensembling. Furthermore, this shows that compared to existing benchmarks, TabArena is closer to correctly representing the currently achievable peak performance.
@@
Tabular Foundation Models Lead on Small Datasets. As most foundation models are currently limited in their applicability, we evaluate them separately in Figure 4, presenting the leaderboard after restricting the datasets to the constraints of TabPFNv2 or TabICL. TabPFNv2 outperforms related approaches by a large margin, establishing tabular foundation models as the go-to solution for datasets within their constraints. Moreover, TabPFNv2 with tuning and post-hoc ensembling again outperforms AutoGluon, confirming the results by Hollmann et al. [5]. We expect tabular foundation models with wider applicability to be released and included in future versions of TabArena.
@@
Figure 4: Leaderboard for TabPFNv2-compatible (left) and TabICL-compatible (right) datasets. For TabPFNv2, we obtain $33$ datasets ($\leq$ 10K training samples, $\leq 500$ features). For TabICL, we obtain $36$ classification datasets ($\leq$ 100K, $\leq 500$). Everything but the datasets is identical to Figure 1.

### v3

@@ Abstract
With the growing popularity of deep learning and foundation models for tabular data, the need for standardized and reliable benchmarks is higher than ever. However, current benchmarks are static. Their design is not updated even if flaws are discovered, model versions are updated, or new models are released. To address this, we introduce TabArena, the first continuously maintained living tabular benchmarking system. To launch TabArena, we manually curate a representative collection of datasets and well-implemented models, conduct a large-scale benchmarking study to initialize a public leaderboard, and assemble a team of experienced maintainers. Our results highlight the influence of validation method and ensembling of hyperparameter configurations to benchmark models at their full potential. While gradient-boosted trees are still strong contenders on practical tabular datasets, we observe that deep learning methods have caught up under larger time budgets with ensembling. At the same time, foundation models excel on smaller datasets. Finally, we show that ensembles across models advance the state-of-the-art in tabular machine learning. We observe that some deep learning models are overrepresented in cross-model ensembles due to validation set overfitting, and we encourage model developers to address this issue. We launch TabArena with a public leaderboard, reproducible code, and maintenance protocols to create a living benchmark available at https://tabarena.ai.
Figure 1: TabArena-v0.1 Leaderboard. We evaluate models under default parameters, tuning, and weighted ensembling [1] of hyperparameters. Since TabICL and TabPFNv2 are not applicable to all datasets, we evaluate them on subsets of the benchmark in Figure˜4 .
@@ 1 Introduction
  • We curate $16$ tabular machine learning models, including $3$ tabular foundation models, and run large-scale experiments in well-tested modeling pipelines used in practice. In total, we trained 25 $000{\sim}$25\,000\,000 instances of models across all our experiments.
@@
TabArena-v0.1 Focus Statement. TabArena focuses on evaluating predictive machine learning models for tabular data. Our long-term vision is to make TabArena representative for all use cases of tabular data. For TabArena-v0.1, we initialize the benchmark focusing on the most common type of tabular machine learning problem: Tabular classification and regression for independent and identically distributed (IID) data, spanning the small to medium data regime. We explicitly leave for future work use cases such as non-IID data (e.g., temporal dependencies, subject groups, or distribution shifts); few-shot predictions, or very small data (e.g., less than 500 training samples); large data (e.g., more than 250 000250\,000 training samples); or other machine learning tasks such as clustering, anomaly detection, subgroup discovery, or survival analysis.
@@ 2 TabArena
TabArena is a living benchmark because of our protocols, which govern the curation of (2.1) models and hyperparameter optimization, (2.2) datasets, and (2.3) evaluation design. Through continuous application and refinement of these protocols, we ensure TabArena remains current and maintained.
@@ 2.1 Models and Hyperparameter Optimization Protocol
TabArena models are powered by three components: (1) implementation in a well-tested modeling framework used in real-world applications; (2) curated hyperparameter optimization protocols; (3) improved validation and ensembling strategies, including ensembling over instances of a single model class. Table˜1 provides an overview of the models benchmarked in TabArena-v0.1.
@@
Linear / Logistic Regression Linear Prior Work + Us [Uncaptioned image]
K-Nearest Neighbors KNN Prior Work + Us [Uncaptioned image]
Implementation Framework.  For implementing models, we rely on functionalities from AutoGluon [19], an established machine learning framework used in practical applications. Each model is implemented within the standardized AbstractModel framework, which aligns with the scikit-learn [24] API, and includes: (1) model-agnostic preprocessing, (2) support for (inner) cross-validation with ensembling, (3) hyperparameter optimization, (4) evaluation metrics, (5) fold-wise training parallelization, (6) (customizable) model-specific preprocessing pipeline, (7) (customizable) early stopping and validation logic, and (8) unit tests. As a result, any model implemented in TabArena can be readily deployed for real-world use cases or within predictive machine learning systems. Moreover, the pipeline logic encompassing models within TabArena is implemented in a tested framework regularly used in real-world applications. Section˜C.1 summarizes further implementation details, and Section˜E.2 includes a detailed protocol for contributing models.
Cross-validation and Post-hoc Ensembles. As can be seen in various Kaggle competitions and academic studies, cf. [25, 8, 26, 20], for most datasets, peak performance requires ensembling strategies. Therefore, we default to using $8$-fold cross-validation (with class-wise stratification for classification) and then employ cross-validation ensembles [27]; which we describe in detail in Section˜C.3. For all foundation models, we refit on training and validation data instead of using cross-validation ensembles, following recommendations from the authors of TabPFN and TabICL. In addition, we evaluate each tunable model using post-hoc ensembling [1] of different hyperparameter configurations, denoted as Tuned + Ensembled; further details are provided in Section˜C.4.
Hyperparameter Optimization.  For each model, we curate a strong hyperparameter search space; for full details, see Appendix C.2. Where possible, we started with the search spaces from the original paper and finalized them in dialogue with the models’ authors. Otherwise, we curated search spaces from prior work. We evaluate $1$ default and a fixed set of $200$ randomly-sampled hyperparameter configurations for all models, except for TabICL and TabDPT. TabICL and TabDPT do not specify hyperparameter optimization (HPO) in the original paper and implementation; thus, we restrict ourselves to evaluating only their default performance. Each hyperparameter configuration is validated using $8$-fold (inner) cross-validation. We use the score of this (inner) cross-validation to select the best hyperparameter configuration.
For practical reasons, we restrict the time to evaluate one configuration on one train split of a dataset to $1$ hour. Our analysis in Section˜A.2 shows that this limit rarely takes effect.
@@ 2.2 Datasets Protocol
Table 2: Comparison of Tabular Benchmarks. We systematically compare prior tabular benchmark studies across six characteristics. Inner and outer splits: the number of splits used for inner or outer validation: 11 for holdout validation; - if the benchmark does not specify; and any other number specifies the total number of splits from (repeated) cross-validation. If a set is given, the benchmark uses different splits for different datasets. Ensembling: Whether the benchmark studies ensembling of configurations for individual models (), uses any other ensembling ((✓)), or uses no ensembling at all (). Manual Curation: Whether the benchmark filters datasets based on criteria beyond simple automation, () or not (). Datasets remaining: the number of datasets remaining after filtering by our criteria. Results available: Whether the benchmark shares no re-usable results (), only metric results ((✓)), or metric results and predictions (). HPO Limit: How hyperparameter optimization was limited in the number of configurations and/or hours.
@@
Many existing benchmarks were curated using semi-automated procedures to collect datasets according to simple characteristics, often to obtain as many datasets as possible. In contrast, we reject the notion of automatically collecting datasets without any sanity check. Instead, we focus on carefully, manually curating a representative collection of datasets. Although some previous benchmarks manually curated data, most of the included datasets do not meet the criteria of IID predictive tabular datasets, as seen in Table˜2. Our work indicates a turning point with the most extensive and conscientious manual curation effort for machine learning on IID tabular data so far.
We define criteria for data selection according to our focus statement and filter 1053 datasets used in 14 prior benchmarks accordingly. Figure 2 describes our selection process. Notably, only the deduplication step and size filters can be automated. Faithfully applying the other criteria requires manual human effort per-dataset, demonstrating the downsides of automated data curation procedures.
@@
Dataset Selection Criteria.  We comprehensively describe our selection criteria in Section˜B.1. In short, we selected datasets that fulfilled the following requirements: (1) The dataset and its predictive machine learning task are unique within our benchmark; (2) The dataset is IID, that is, a random split is appropriate for the underlying original task; (3) The dataset is not from a non-tabular modality, such as images, where it is unclear whether tabular machine learning is a reasonable alternative to domain-specific methods; (4) The dataset stems from a real random distribution, and is not generated, e.g., from a deterministic function; (5) The dataset was published explicitly for a predictive modeling task in a real-world application; (6) The dataset is small-to-medium-sized, i.e., it has at least $500$ and at most 250 000250\,000 train samples; (7) We can use a version of the dataset without pre-applied problematic preprocessing, such as irreversible data leaks; (8) The dataset was originally published with a license allowing for scientific usage; (9) The dataset and its structured metadata can be automatically downloaded via a public API, or we are allowed to upload the dataset to a public API; (10) The dataset and its predictive task do not raise (obvious) ethical concerns.
@@
Figure˜3 summarizes dataset characteristics; we share per-dataset details and domain coverage in Section˜B.2. Throughout our curation process, we noticed two trends. First, across benchmarks, the versions of datasets were often inconsistent. Benchmarks used the same name and source for a dataset, but with different preprocessing, features, or even targets. Thus, dataset-specific performance comparisons across benchmarks were often invalid. Second, the number of datasets that are truly suitable for benchmarking IID tabular data approaches is surprisingly small after carefully inspecting the tasks represented by datasets. We share more noteworthy observations in Section˜B.3.
Call for Data Contributions.TabArena is a living benchmark, and the datasets we curated are a significant part of this living, continuously maintained system. As mentioned above, we invite the community to scrutinize our curation efforts. Moreover, to keep TabArena up-to-date, we implore the broader scientific community to share their tabular data publicly. We are also actively looking for contributions of data outside of TabArena-v0.1’s focus, i.e., data that is non-IID, tiny, or large, as well as tasks where extracing features from other modalities is still considered a reasonable approach. Likewise, we invite the community to curate additional existing datasets. We provide a detailed protocol to contribute a new dataset to TabArena in Section˜E.3.
@@ 2.3 Evaluation Design Protocol
The TabArena leaderboard aims to assess the state-of-the-art for predictive machine learning on tabular data. With this in mind, we design the evaluation to produce a reliable and representative leaderboard. To guard against randomness in the data and method significantly impacting our conclusions, we repeat our experiments per dataset. We detail the various sources of randomness in TabArena in Section˜D.2. We employ a dataset-specific repetition strategy using more repeats for smaller datasets and fewer for larger datasets, because the significance of randomness decreases with dataset size. The strategy is as follows: (I) for datasets with less than 2 5002\,500 samples, we use $10$ times repeated $3$-fold outer cross-validation; (II) for all other datasets, we use $3$ repeats. For classification tasks, we use class-wise stratified cross-validation.
Evaluation Metric. We evaluate models using the Elo rating system [38]. Elo is a pairwise comparison-based rating system where each model’s rating predicts its expected win probability against others, with a 400-point Elo gap corresponding to a 10 to 1 (91%) expected win rate. We calibrate 1000 Elo to the performance of our default random forest configuration across all figures, and perform 200 rounds of bootstrapping to obtain 95% confidence intervals, similar to [39], see Section˜D.1. In our main results, Elo scores are computed using ROC AUC for binary classification, log-loss for multiclass classification, and RMSE for regression.
We aim to provide a ranking of models as part of TabArena. Every aggregation metric used for ranking has its own pitfalls, and none is perfect; however, we argue that Elo aligns most closely with our goals and has therefore been selected as our primary evaluation metric. Elo is based on pairwise comparison scoring, which only considers wins, ties, or losses and neglects the magnitude of performance differences. This can be a disadvantage, as minor performance differences may be considered irrelevant by practitioners in some applications. At the same time, this means that each dataset contributes equally to Elo, hence the aggregation is not biased towards certain domains or dataset properties (e.g., small or non-noisy datasets). For TabArena, this is a key advantage because we want to create a benchmark whose results are representative of all domains and datasets.
While we use Elo as our primary evaluation metric, we also track and present additional aggregation metrics in Section˜3 and on the leaderboard. Users can re-rank the leaderboard according to alternative metrics, such as Improvability. Improvability measures the performance of methods relative to the best method, and is therefore sensitive to the magnitude of performance differences; see Section˜A.1 for details. We also share scripts to generate the leaderboard, evaluation plots, and inspect the results.
TabArena Reference Pipeline. Following the recommendations of [11], we include a reference pipeline in our benchmark. This reference pipeline is applied independently of the tuning protocol and constraints we constructed for models within TabArena. It aims to represent the performance easily achievable by a practitioner on a dataset.
@@
Living Benchmark.TabArena marks the start of an open-source initiative towards a continuously updated, collectively shared assessment of the state-of-the-art in tabular machine learning. Therefore, as one of the most crucial parts of TabArena, we define a protocol for researchers and practitioners to submit models to the live leaderboard. The protocol is detailed in Section˜E.4
@@ 3 Results
To initialize the leaderboard, we run $16$ curated models (Section˜2.1) on $51$ curated datasets (Section˜2.2) within TabArena’s evaluation design (Section˜2.3). Likewise, we evaluate AutoGluon, the reference pipeline, on all datasets. We run TabM, ModernNCA, and the foundation models on GPU and all other models, as well as AutoGluon, on CPU. We give all runs $32$ GB of RAM and $100$ GB of disk space. We perform CPU runs in the cloud via Amazon Web Services on an M6i.2xlarge EC2 instance (eight cores Intel Xeon CPU). We perform GPU runs on an NVIDIA L40S with 48 GB VRAM and eight cores of an AMD EPYC 9334 CPU. Based on our logs, the compute time for all our experiments without parallelization, including overhead such as scheduling, is ${\sim}15$ wall-clock years.
@@
Figure˜1 shows our main results, the leaderboard of TabArena-v0.1, which includes the performance of methods with default parameters, a single configuration on validation data after hyperparameter tuning, and weighted post-hoc ensembling [1] of different hyperparameter configurations. Furthermore, we present one leaderboard encompassing all models in Section˜A.1, where we impute missing datasets, representing the live leaderboard available on tabarena.ai. To provide a wide range of perspectives, we also share additional versions of the leaderboard in Section˜A.1, such as using alternative metrics or task-wise dataset subsets. Finally, we analyze the statistical significance of our comparison in Section˜A.5.
@@
Post-Hoc Ensembled Deep Learning Models Dominate the Leaderboard. In line with previous work [33], CatBoost is ranked first in the conventional tuning regime (Figure˜1). However, the TabArena-v0.1 leaderboard reveals that after post-hoc ensembling, neural networks are the strongest single models on average in TabArena-v0.1. Our results show that the peak performance of models is misrepresented unless post-hoc ensembling is used. This is evident in the observation that the top three models in our leaderboard (TabM, LightGBM, RealMLP; see Figure˜1) would all be worse than the actual fourth-best model (CatBoost) without post-hoc ensembling. While practitioners must trade off the increased predictive performance with the increased inference cost (see Figure˜6) [40], peak performance requires post-hoc ensembling. Furthermore, this shows that compared to existing benchmarks, TabArena is closer to correctly representing the currently achievable peak performance.
@@
Tabular Foundation Models Lead on Small Datasets. As most foundation models are currently limited in their applicability, we evaluate them separately in Figure˜4, presenting the leaderboard after restricting the datasets to the constraints of TabPFNv2 or TabICL. TabPFNv2 outperforms related approaches by a large margin, establishing tabular foundation models as the go-to solution for datasets within their constraints. Moreover, TabPFNv2 with tuning and post-hoc ensembling again outperforms AutoGluon, confirming the results by Hollmann et al. [5]. We expect tabular foundation models with wider applicability to be released and included in future versions of TabArena.
@@
Figure 4: Leaderboard for TabPFNv2-compatible (left) and TabICL-compatible (right) datasets. For TabPFNv2, we obtain $33$ datasets ($\leq$ 10K training samples, $\leq 500$ features). For TabICL, we obtain $36$ classification datasets ($\leq$ 100K, $\leq 500$). Everything but the datasets is identical to Figure˜1.
Pareto Fronts and Tuning Trajectories Reveal Efficiency Tradeoffs. The Pareto front of improvability and inference time in Figure˜5 (left) reveals that tuned EBM (EBM-T) and tuned CatBoost (CatBoost-T) shine at inference time. The next Pareto points with strong improvements come with an increase in inference time of $\sim 15\times$ for TabM-TE and $\sim 100\times$ for RealMLP-TE. Figure˜5 (right) shows that gradient-boosted trees have strong performance given their training cost. RealMLP only starts to dominate them after a considerable amount of training time with an ensemble of 25+ configurations. For additional discussion of the tuning trajectories, see Section˜A.6.
Figure 5: (Left) Pareto front of improvability and inference time. We report the median inference time per 1000 samples across all datasets. (Right) Improvability tuning trajectories. Time is shown as the tuning time with points from left to right marking ensembles of increasing numbers of random configurations (1, 2, 5, 10, 25, 50, 100, 150, 201). The trajectories are sampled 20 times from all trials and averaged. The right-most highlighted points use all configurations.